# HEIE: MLLM-Based Hierarchical Explainable AIGC Image Implausibility Evaluator

### v1

@@ Abstract
AIGC images are prevalent across various fields, yet they frequently suffer from quality issues like artifacts and unnatural textures. Specialized models aim to predict defect region heatmaps but face two primary challenges: (1) lack of explainability, failing to provide reasons and analyses for subtle defects, and (2) inability to leverage common sense and logical reasoning, leading to poor generalization. Multimodal large language models (MLLMs) promise better comprehension and reasoning but face their own challenges: (1) difficulty in fine-grained defect localization due to the limitations in capturing tiny details; and (2) constraints in providing pixel-wise outputs necessary for precise heatmap generation. To address these challenges, we propose HEIE: a novel MLLM-Based Hierarchical Explainable image Implausibility Evaluator. We introduce the CoT-Driven Explainable Trinity Evaluator, which integrates heatmaps, scores, and explanation outputs, using CoT to decompose complex tasks into subtasks of increasing difficulty and enhance interpretability. Our Adaptive Hierarchical Implausibility Mapper synergizes low-level image features with high-level mapper tokens from LLMs, enabling precise local-to-global hierarchical heatmap predictions through an uncertainty-based adaptive token approach. Moreover, we propose a new dataset: Expl-AIGI-Eval, designed to facilitate interpretable implausibility evaluation of AIGC images. Our method demonstrates state-of-the-art performance through extensive experiments.
@@ 1 Introduction
Recently, Artificial Intelligence Generated Content (AIGC) has rapidly advanced, leading to the widespread use of AIGC images across various fields. However, they often face quality issues such as artifacts, unnatural textures, and problematic regions. Accordingly, many studies focus on the evaluation of AIGC images [44, 42, 19, 26]. Current evaluations primarily yield scalar scores, simple 0-1 values, failing to identify specific defect regions, which hinders targeted improvements in generative models.
@@
  1. 1.
    Lack of explainability. As shown in Fig. 1 (a), implausibility regions in AIGC images can be complex, such as twisted fingers or overly large eye corners. Specialized small models fail to provide detailed explanations for these subtle defects, making it difficult for general users to understand and optimize the AIGC models.
  2. 2.
    Lack of general knowledge. These small specialized models are trained on limited datasets and lack comprehensive world knowledge, common sense, and logical reasoning skills. As a result, they struggle to generalize to out-of-domain samples during practical usage.
@@
  1. 1.
    Fine-Grained Defect Localization: While MLLMs excel in capturing global image structure, they struggle with pinpointing fine-grained defects. Tiny defect areas like the corners of eyes or eyebrows (Fig. 1 (b)) are often neglected by strong MLLMs such as GPT-4o.
  2. 2.
    Pixel-Wise Output Limitations: General-purpose MLLMs handle various modalities (text, images, videos, etc.) but usually output text only, lacking the capability to produce pixel-level defect heatmaps (Fig. 1 (b)).
@@
    We construct a high-quality explainable AIGC image implausibility evaluation dataset, which will be publicly available.
Our method achieves state-of-the-art performance on various datasets and tasks. Extensive comparisons and ablation studies validate the effectiveness of our approach.
@@ 2.2 AIGC Image Evaluation
Effective evaluation are required to optimize generative models. Many studies have released datasets to emulate human assessment [44, 42, 19, 14, 21]. Large-scale visual understanding models are now commonly used for reliable evaluation due to their strong semantic capabilities [28, 41, 50]. Prior research has shown that MLLMs are proficient in both low-level visual evaluations [37, 39] and high-level semantic comprehension [52, 49]. Some studies  [38, 34, 17, 40] activate the visual evaluation capabilities of multimodal MLLMs through data fine-tuning.
@@
Recently, HumanRefiner [10] and RichHF  [23] have introduced tasks focused on detecting fine-grained problematic regions in images. HumanRefiner targets defects in AIGC-generated human images, while RichHF encompasses defects in various objects. However, these approaches rely on proprietary small models to detect defect regions. These models are limited in common sense and knowledge, resulting in insufficient generalization capabilities in complex scenarios. Additionally, proprietary small models lack interpretability and cannot provide detailed reasons or explanations  [2, 9] for the detected issues.
@@ 3.2.1 MLLM-Based Implausibility Mapper
To obtain fine-grained, precise image implausibility regions, pixel-level mask outputs are necessary. However, most MLLMs only provide text information and cannot produce heatmap predictions. To address this, we design the Image Implausibility Mapper module for heatmap prediction. We introduced a special map token [MAP] to extract implausibility region information from LLM. And we use image features $F$ from the image encoder and [MAP] token features $T$ from the LLM for a two-way cross-attention calculation [18, 1, 35], as shown in Eq. LABEL:2waytrans. To fuse the [MAP] token embeddings with image features, each input alternates as the query and value during cross-attention operations. Specifically, The special map token embeddings $T$ and image features $F$ server as $E_{\text{text}}$ and $E_{\text{img}}$ respectively in the following two-way cross-attention:
$C_{\text{text\textrightarrow img}}=\text{CrossAtten}(\text{SelfAtten}(E_{\text{text}}),E_{\text{img}},E_{\text{img}})$ (1)
$E_{\text{text}^{{}^{\prime}}}=\text{mlp}(\text{norm}(C_{\text{text\textrightarrow img}}))$
$C_{\text{img\textrightarrow text}}=\text{CrossAtten}(E_{\text{img}},E_{\text{text}^{{}^{\prime}}},E_{\text{text}^{{}^{\prime}}})$
$O_{2way\_CA}=\text{norm}(C_{\text{img\textrightarrow text}})$
By stacking the above two-way cross-attention twice, we get the predicted heatmap $H$.
@@ 3.3.2 CoT-Driven Explainable System
1. Image Description: The LLM describes the image, capturing key elements in natural language.
2. Problematic Regions Identification: Based on the description and the understanding of the image, the LLM identifies potential problematic areas for further analysis.
3. Special map token: Based on the preceding Problematic Regions Identification, the LLM injects defect region information into the special map token for use by the implausibility mapper.
4. Analysis of Issues: Based on the preceding implausibility localization, the LLM provides detailed text explanations for each region, including type, causes, and so on.
5. Special Score Token: Based on the comprehensive understanding, the LLM incorporates overall information into a special score token for use by our Verisimilitude Scorer.
@@ 3.4 Expl-AIGI-Eval Dataset
Existing AIGC image implausibility datasets, such as RichHF-18K [23] and AbHuman [10], include AIGC images and annotated defect regions but lack interpretability explanations. To address this, we introduce Expl-AIGI-Eval: Explainable AIGC Image Implausibility Evaluation. We provide detailed analyses and explanations for each implausibility region. The annotation process is conducted in three stages, shown in Fig. 3.
1. Visual Prompting: Research [6] indicates that LLMs effectively adapt to visual prompts. Consequently, we annotate defect regions directly on images, ensuring visibility while preserving the image’s integrity and facilitating subsequent use by Claude-3.5-sonnet.
2. LLM Free-Form Outputting: Using Claude-3.5-sonnet, we generate detailed analyses from visual prompts. Prior studies [33] reveal that fixed-format enforcement degrades performance, so we allow free-form outputs for defect causes and analyses.
3. In-Context Learning-Based Formatting: To ensure consistency, we employ in-context learning to guide the GPT-4o in refining and formatting information.
@@
In summary, our Expl-AIGI-Eval dataset comprises 42,760 AIGC images divided into two parts: RichHF-18K [23] and AbHuman [10]. We create 20 diverse query expressions per image for LLM instructions. We manually review 20% of the generated content for quality assurance. Sensitive images related to pornography, privacy, animal protection, and political topics are filtered out to avoid ethical issues.
@@ 4.1.2 Evaluation Metrics
We follow the method in [23] for evaluation. We report the MSE for all samples, including images with no problematic region. For images with implausibility regions, we provide evaluation metrics for the saliency heatmaps, such as KLD, AUC-Judd, SIM, and CC, as referenced in [5]. For the score prediction tasks, we report the Pearson linear correlation coefficient (PLCC) and Spearman rank correlation coefficient (SRCC) [5]. Additionally, since the RAHF model [23] is not open source and its training is complex, we can only provide the results reported in the RichHF paper and cannot test the actual performance of the RAHF model.
@@ 4.2 Comparison with State-of-the-art
We evaluate our method against the RichHF-18K benchmark [23], as presented in Tab. 1. Conventional vision models typically lack physical common sense and struggle with high-level semantic understanding, complicating their ability to perceive and recognize content that goes against human intuition. In contrast, our method excels in accurately predicting heatmaps. Given the strong correlation between heatmap accuracy and image plausibility scores, our approach provides quantitative metrics that align with human evaluations. In comparison, our model achieved SOTA performance.
As shown in Tab. 2, we evaluate the performance of image implausibility explanations on our Expl-AIGI-Eval Dataset. We employe three evaluation methods to evaluate the models’ text explanation capability: (1) Compare the perplexity (ppl) of model outputs to ground truth; (2) Employ GPT-4o to compare the content similarity between the model outputs and the ground truth. (3) Conduct human evaluations of the model-generated explanations. Our carefully designed CoT approach significantly enhances our analysis and explanation capabilities for AIGC image implausibility. Our textual analysis outperforms both other open-source MLLMs and surpasses the general-purpose SOTA models like GPT-4o and Claude-3.5-sonnet. Across various textual explanation metrics, our model achieves SOTA performance.
Tab. 3 shows the results on the AbHuman dataset [10], which focuses more on the flaws in AI-generated images of humans. Baseline models underperform due to limited common sense and reasoning skills. In contrast, our model achieves SOTA results. Additionally, the RAHF model’s performance remains uncertain due to undisclosed results and its complex, closed-source training process.
@@ 4.3 Zero-shot Domain Generalization Results
Tab. 4 presents zero-shot cross-domain generalization experiments. We train on either the RichHF-18K or AbHuman datasets and perform zero-shot inference on the other. Cross-domain tasks [48, 13], common in real-world applications [46, 47], test a model’s fundamental world knowledge beyond domain-specific fitting [12]. As shown, smaller models suffer significant accuracy drops, losing defect assessment ability. In contrast, our model excels, leveraging MLLMs’ strong common-sense knowledge and zero-shot capabilities.
@@ 4.4 Ablation Study
Ablation Study for CoT-Driven Explainable System. As shown in Tab. 7, we evaluate our CoT Instruction. The first experiment without CoT Text (w/o CoT Text) reports results lacking our CoT Instruction, while the second (w/ CoT Text) shows outcomes with it. Comparing these, CoT Prompt significantly enhances heatmap and score predictions, demonstrating its effectiveness in guiding logical reasoning in LLMs. The third experiment (w/ GT CoT Text) uses ground truth CoT text labels, including image descriptions and problematic region identification, as instruction. These further boost prediction accuracy, affirming CoT’s value in enhancing LLM [45, 43] reasoning.

### v2

@@ Abstract
AIGC images are prevalent across various fields, yet they frequently suffer from quality issues like artifacts and unnatural textures. Specialized models aim to predict defect region heatmaps but face two primary challenges: (1) lack of explainability, failing to provide reasons and analyses for subtle defects, and (2) inability to leverage common sense and logical reasoning, leading to poor generalization. Multimodal large language models (MLLMs) promise better comprehension and reasoning but face their own challenges: (1) difficulty in fine-grained defect localization due to the limitations in capturing tiny details, and (2) constraints in providing pixel-wise outputs necessary for precise heatmap generation. To address these challenges, we propose HEIE: a novel MLLM-Based Hierarchical Explainable Image Implausibility Evaluator. We introduce the CoT-Driven Explainable Trinity Evaluator, which integrates heatmaps, scores, and explanation outputs, using CoT to decompose complex tasks into subtasks of increasing difficulty and enhance interpretability. Our Adaptive Hierarchical Implausibility Mapper synergizes low-level image features with high-level mapper tokens from LLMs, enabling precise local-to-global hierarchical heatmap predictions through an uncertainty-based adaptive token approach. Moreover, we propose a new dataset: Expl-AIGI-Eval, designed to facilitate interpretable implausibility evaluation of AIGC images. Our method demonstrates state-of-the-art performance through extensive experiments. Our project is at https://yfthu.github.io/HEIE/.
@@ 1 Introduction
Recently, Artificial Intelligence Generated Content (AIGC) has rapidly advanced, leading to the widespread use of AIGC images across various fields. However, they often face quality issues such as artifacts, unnatural textures, and problematic regions. Accordingly, many studies focus on the evaluation of AIGC images [51, 50, 22]. Current evaluations primarily yield scalar scores, simple 0-1 values, failing to identify specific defect regions, which hinders targeted improvements in generative models.
@@
(1) Lack of explainability. As shown in Fig. 1 (a), implausibility regions in AIGC images can be complex, such as twisted fingers or overly large eye corners. Specialized small models fail to provide detailed explanations for these subtle defects, making it difficult for general users to understand and optimize the AIGC models.
(2) Lack of general knowledge. These small specialized models are trained on limited datasets and lack comprehensive world knowledge, common sense, and logical reasoning skills. As a result, they struggle to generalize to out-of-domain samples during practical usage.
@@
(1) Fine-Grained Defect Localization: While MLLMs excel in capturing global image structure, they struggle with pinpointing fine-grained defects. Tiny defect areas like the corners of eyes or eyebrows (Fig. 1 (b)) are often neglected by strong MLLMs such as GPT-4o.
(2) Pixel-Wise Output Limitations: General-purpose MLLMs handle various modalities (text, images, videos, etc.) but usually output text only, lacking the capability to produce pixel-level defect heatmaps (Fig. 1 (b)).
@@
    We construct and publish a high-quality explainable AIGC image implausibility evaluation dataset. Our method achieves state-of-the-art performance on various datasets and tasks. Extensive comparisons and ablation studies validate the effectiveness of our approach.
@@ 2.2 AIGC Image Evaluation
Effective evaluation are required to optimize generative models. Many studies have released datasets to emulate human assessment [51, 50, 22, 17, 24]. Large-scale visual understanding models are now commonly used for reliable evaluation due to their strong semantic capabilities [30, 49, 56]. Prior research has shown that MLLMs are proficient in both low-level visual evaluations [45, 47] and high-level semantic comprehension [58, 55, 36, 35, 34]. Some studies  [46, 39, 20, 48] activate the visual evaluation capabilities of multimodal MLLMs through data fine-tuning.
@@
Recently, HumanRefiner [13] and RichHF  [26] have introduced tasks focused on detecting fine-grained problematic regions in images. HumanRefiner targets defects in AIGC-generated human images, while RichHF encompasses defects in various objects. However, these approaches rely on specialized small models to detect defect regions. These models are limited in common sense and knowledge, resulting in insufficient generalization capabilities in complex scenarios. Additionally, specialized small models lack interpretability and cannot provide detailed reasons or explanations  [2, 12] for the detected issues.
@@ 3.2.1 MLLM-Based Implausibility Mapper
To obtain fine-grained, precise image implausibility regions, pixel-level mask outputs are necessary. However, most MLLMs only provide text information and cannot produce heatmap predictions. To address this, we design the Image Implausibility Mapper module for heatmap prediction. We introduced a special map token [MAP] to extract implausibility region information from LLM. In LLM, the [MAP] token generates many feature layers, and its final linear layer feature is used to predict the next token for LLM. We select this final LLM layer’s hidden state of the [MAP] token as the features $T$. Then we use image features $F$ from the image encoder and [MAP] token features $T$ from the LLM for a two-way cross-attention calculation [21, 1, 40], as shown in Eq. LABEL:2waytrans.
$C_{\text{T\textrightarrow F}}=\text{CrossAtten}(\text{SelfAtten}(T),F,F)$ (1)
$T^{\prime}=\text{mlp}(\text{norm}(C_{\text{T\textrightarrow F}}))$
$C_{\text{F\textrightarrow T}}=\text{CrossAtten}(F,T^{\prime},T^{\prime})$
$O_{2way\_CA}=\text{norm}(C_{\text{F\textrightarrow T}})$
To fuse the [MAP] token features $T$ with image features $F$, each input alternates as the query and value during cross-attention operations. By stacking the above two-way cross-attention Eq. (LABEL:2waytrans) twice, we get the predicted heatmap $H$.
@@ 3.3.2 CoT-Driven Explainable System
(1) Image Description: The LLM describes the image, capturing key elements in natural language.
(2) Problematic Regions Identification: Based on the description and the understanding of the image, the LLM identifies potential problematic areas for further analysis.
(3) Special map token: Based on the preceding Problematic Regions Identification, the LLM injects defect region information into the special map token for use by the implausibility mapper.
(4) Analysis of Issues: Based on the preceding implausibility localization, the LLM provides detailed text explanations for each region, including type, causes, and so on.
(5) Special Score Token: Based on the comprehensive understanding, the LLM incorporates overall information into a special score token for use by our Verisimilitude Scorer.
@@ 3.4 Expl-AIGI-Eval Dataset
Existing AIGC image implausibility datasets, such as RichHF-18K [26] and AbHuman [13], include AIGC images and annotated defect regions but lack interpretability explanations. To address this, we introduce Expl-AIGI-Eval: Explainable AIGC Image Implausibility Evaluation. We provide detailed analyses and explanations for each implausibility region. The annotation process is conducted in the following three stages, shown in Fig. 3.
(1) Visual Prompting: We follow [7] to add visual prompts. Specifically, the original dataset provides masks of the defective regions for each image. We draw a red contour for each defective region, clearly highlighting these areas, as shown in Stage 1 in Fig. 3.
(2) LLM Free-Form Outputting: Using Claude-3.5-sonnet [3], we generate detailed analyses of the defect regions. Previous study [38] indicates that enforcing fixed-format outputs impairs the performance of LLM, so we allow the LLM to produce free-form outputs for analyzing defect causes.
(3) In-Context Learning-Based Formatting: As shown in Stage-3 (Fig.3), we employ GPT-4o and in-context learning to format the defect analysis texts obtained in Stage 2.
@@
Sensitive images related to pornography, privacy, animal protection, and political topics are filtered out to avoid ethical issues.
@@ 4.1.2 Evaluation Metrics
We follow the method in [26] for evaluation. We report the MSE for all samples, including images with no problematic region. For images with implausibility regions, we provide evaluation metrics for the saliency heatmaps, such as KLD, AUC-Judd, SIM, and CC, as referenced in [6]. For the score prediction tasks, we report the Pearson linear correlation coefficient (PLCC) and Spearman rank correlation coefficient (SRCC) [6]. Additionally, since the RAHF model [26] is not open source, we provide its results reported in its paper.
@@ 4.2 Comparison with State-of-the-art
We evaluate our method on the RichHF-18K benchmark [26], as presented in Tab. 1. Conventional vision models typically lack physical common sense and struggle with high-level semantic understanding, complicating their ability to perceive and recognize content that goes against human intuition. In contrast, our method excels in accurately predicting heatmaps. Given the strong correlation between heatmap accuracy and image plausibility scores, our approach provides quantitative metrics that align with human evaluations. In comparison, our model achieved SOTA performance.
As shown in Tab. 2, we evaluate the performance of image implausibility explanations on our Expl-AIGI-Eval Dataset. We employ three evaluation methods to evaluate the models’ text explanation capability: (1) Compare the perplexity (ppl) of model outputs to ground truth; (2) Employ GPT-4o to compare the content similarity between the model outputs and the ground truth. (3) Conduct human evaluations of the model-generated explanations. Our carefully designed CoT approach significantly enhances our analysis and explanation capabilities for AIGC image implausibility. Our textual analysis outperforms both other open-source MLLMs and surpasses the general-purpose SOTA models like GPT-4o and Claude-3.5-sonnet. Across various textual explanation metrics, our model achieves SOTA performance.
Tab. 3 shows the results on the AbHuman dataset [13], which focuses more on the flaws in AI-generated images of humans. Baseline models underperform due to limited common sense and reasoning skills. In contrast, our model achieves SOTA results.
@@ 4.3 Zero-shot Domain Generalization Results
Tab. 4 presents zero-shot cross-domain generalization experiments. We train on either the RichHF-18K or AbHuman datasets and perform zero-shot inference on the other. Cross-domain tasks [54, 16], common in real-world applications [42, 41, 11, 10, 53, 52], test a model’s fundamental world knowledge beyond domain-specific fitting [15, 43]. As shown, smaller models suffer significant accuracy drops, losing defect assessment ability. In contrast, our model excels, leveraging MLLMs’ strong common-sense knowledge and zero-shot capabilities.
@@ 4.4 Ablation Study
Ablation Study for CoT-Driven Explainable System. As shown in Tab. 7, we evaluate our CoT Instruction. The first experiment without CoT Text (w/o CoT Text) reports results lacking our CoT Instruction, while the second (w/ CoT Text) shows outcomes with it. Comparing these, CoT Prompt significantly enhances heatmap and score predictions, demonstrating its effectiveness in guiding logical reasoning in LLMs. The third experiment (w/ GT CoT Text) uses ground truth CoT text labels, including image descriptions and problematic region identification, as instruction. These further boost prediction accuracy, affirming CoT’s value in enhancing LLM reasoning.